OpenStudy (anonymous):

how do i factor a^2-b^2-4b-4?

OpenStudy (steve816):

I don't think it's factorable...

OpenStudy (steve816):

\[a^2-b^2-4b-4\]

OpenStudy (anonymous):

Oh...I got stuck at (a+b)(a-b)-4(b+1). Thanks! For some reason, if you insert the equation into http://www.wolframalpha.com/widgets/view.jsp?id=e7040eab64a24724666a4991fb077bd2, it can be factorable

OpenStudy (anonymous):

^^wolfram factoring calculator http://www.wolframalpha.com/widgets/view.jsp?id=e7040eab64a24724666a4991fb077bd2

Directrix (directrix):

a^2 - b^2 -4b - 4 = a^2 - ( b^2 +4b + 4)

Directrix (directrix):

This expression is the difference of the perfect square a^2 and the perfect square trinomial ( b^2 +4b + 4)

OpenStudy (anonymous):

Yes, that was what the answer was suppose to be. I always end up grouping the wrong polynomials. Thanks.

Directrix (directrix):

a^2 - ( b^2 + 4b + 4) = a^2 - [(b + 2)*(b+2)] = a^2 - ( b + 2)^2 =

Directrix (directrix):

a^2 - ( b + 2)^2 = ? @one234 (I grouped the wrong ones together on my first try, too. :) )

OpenStudy (anonymous):

(a+b+2)(a-b+2)

Directrix (directrix):

a^2 - ( b + 2)^2 = [ a + (b + 2 ) ] * [ a - ( b + 2 ) ]

Directrix (directrix):

I think you have a sign error in distributing the negative here: [ a - ( b + 2 ) ]

Directrix (directrix):

- ( b + 2 ) The negative applies to both the b and the 2. So, what would be the correct answer? @one2345 (Or, did I mess up?)

OpenStudy (anonymous):

(a-b-2) I didn't realize it was necessary to use the parentheses. :P

Directrix (directrix):

[ a + (b + 2 ) ] * [ a - ( b + 2 ) ] = (a + b + 2) * (a - b - 2)

OpenStudy (anonymous):

Thank you again.

OpenStudy (anonymous):

would you mind showing me how to do one more factoring problem?

Directrix (directrix):

Glad to help. Try the problem I attached. Post what you get, okay?

OpenStudy (anonymous):